OpenStudy (anonymous):

g(x)= 4/x-4, Find g(x-1)

mathslover (mathslover):

Is it : \(g(x) = \cfrac{4}{x-4}\) or \(g(x) = \cfrac{4}{x} - 4\) ?

OpenStudy (anonymous):

it is the first one

mathslover (mathslover):

Okay, so here is it what you have to do : On the place of x , plug-in : x-1 and simplify. That is : \(g(x-1) = \cfrac{4}{(x-1)-4}\)

OpenStudy (anonymous):

that is what i thought THANKS

mathslover (mathslover):

Good work and you're welcome Also, \(\LARGE \color{blue}{\mathsf{Welcome}} \space \color{orange}{\mathbb{To}} \space \color{red}{\textbf{OpenStudy!!!}}\)

OpenStudy (anonymous):

thanks i always wanted to study and be online at the same time